# Interplay of Zeeman field, Rashba spin-orbit interaction, and superconductivity: Transition temperature and quasiparticle excitations

### v1

@@ Appendix G Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC
$H_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+g|\mathbf{k}|\cos\theta_{\mathbf{k}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}e^{-i\varphi_{\mathbf{k}}}&[-d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]e^{i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})e^{i\eta_{\mathbf{k}}}\\ g|\mathbf{k}|\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}&\xi_{\mathbf{k}}-g|\mathbf{k}|\cos\theta_{\mathbf{k}}&d^{\prime}_{z}(\mathbf{k})e^{i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]e^{i\eta_{\mathbf{k}}}\\ [-d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]e^{-i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})e^{-i\eta_{\mathbf{k}}}&-\xi_{\mathbf{k}}+g|\mathbf{k}|\cos\theta_{\mathbf{k}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}\\ d^{\prime}_{z}(\mathbf{k})e^{-i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]e^{-i\eta_{\mathbf{k}}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}e^{-i\varphi_{\mathbf{k}}}&-\xi_{\mathbf{k}}-g|\mathbf{k}|\cos\theta_{\mathbf{k}}\end{pmatrix},$ (89)
@@
$C_{\mathbf{k}}$ $=U_{1\mathbf{k}}C^{\prime}_{\mathbf{k}}=\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}&0&0\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&0&0\\ 0&0&\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ 0&0&\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}\end{pmatrix}\begin{pmatrix}c_{\mathbf{k}+}\\ c_{\mathbf{k}-}\\ c_{-\mathbf{k}-}^{\dagger}\\ c_{-\mathbf{k}+}^{\dagger}\end{pmatrix},$ (90)
@@
$H^{\prime}_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+g|\mathbf{k}|&0&[-d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]e^{i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})e^{i\gamma_{\mathbf{k}}}\\ 0&\xi_{\mathbf{k}}-g|\mathbf{k}|&d_{3}(\mathbf{k})e^{i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]e^{i\gamma_{\mathbf{k}}}\\ [-d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]e^{-i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})e^{-i\gamma_{\mathbf{k}}}&-\xi_{\mathbf{k}}+g|\mathbf{k}|&0\\ d_{3}(\mathbf{k})e^{-i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]e^{-i\gamma_{\mathbf{k}}}&0&-\xi_{\mathbf{k}}-g|\mathbf{k}|\end{pmatrix},$ (91)
@@
$=\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}&0&0\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&0&0\\ 0&0&\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ 0&0&\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}\end{pmatrix}\cdot\frac{1}{2\sqrt{E_{\mathbf{k}\perp}}}$
@@
$\begin{pmatrix}e^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&e^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-e^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-e^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ \sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\end{pmatrix},$
@@
$d_{12}(\mathbf{k})e^{i\Omega_{\mathbf{k}}}$ $=d_{1}(\mathbf{k})+id_{2}(\mathbf{k}),$
@@
$u_{\mathbf{k}}$ $=\frac{e^{i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}\cos\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}-e^{i\Omega_{\mathbf{k}}}\sin\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}2+}b_{\mathbf{k}-}&\cos\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}+e^{i\Omega_{\mathbf{k}}}\sin\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}1-}b_{\mathbf{k}+}\\ e^{i\varphi_{\mathbf{k}}}(\sin\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}+e^{i\Omega_{\mathbf{k}}}\cos\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}2+}b_{\mathbf{k}-})&e^{i\varphi_{\mathbf{k}}}(\sin\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}-e^{i\Omega_{\mathbf{k}}}\cos\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}1-}b_{\mathbf{k}+})\end{pmatrix},$ (98)
$v_{\mathbf{k}}$ $=\frac{1}{2}\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}+e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}&\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}-e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}\\ e^{i\varphi_{\mathbf{k}}}(\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}-e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-})&e^{i\varphi_{\mathbf{k}}}(\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}+e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+})\end{pmatrix},$

### v2

@@ Appendix H Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC
$H_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+g|\mathbf{k}|\cos\theta_{\mathbf{k}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&[-d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}}\\ g|\mathbf{k}|\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&\xi_{\mathbf{k}}-g|\mathbf{k}|\cos\theta_{\mathbf{k}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{i\eta_{\mathbf{k}}}\\ [-d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{-i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{-i\eta_{\mathbf{k}}}&-\xi_{\mathbf{k}}+g|\mathbf{k}|\cos\theta_{\mathbf{k}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\\ d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{-i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{-i\eta_{\mathbf{k}}}&g|\mathbf{k}|\sin\theta_{\mathbf{k}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&-\xi_{\mathbf{k}}-g|\mathbf{k}|\cos\theta_{\mathbf{k}}\end{pmatrix},$ (89)
@@
$C_{\mathbf{k}}$ $=U_{1\mathbf{k}}C^{\prime}_{\mathbf{k}}=\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}&0&0\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&0&0\\ 0&0&\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ 0&0&\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\end{pmatrix}\begin{pmatrix}c_{\mathbf{k}+}\\ c_{\mathbf{k}-}\\ c_{-\mathbf{k}-}^{\dagger}\\ c_{-\mathbf{k}+}^{\dagger}\end{pmatrix},$ (90)
@@
$H^{\prime}_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+g|\mathbf{k}|&0&[-d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]\mathrm{e}^{i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})\mathrm{e}^{i\gamma_{\mathbf{k}}}\\ 0&\xi_{\mathbf{k}}-g|\mathbf{k}|&d_{3}(\mathbf{k})\mathrm{e}^{i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]\mathrm{e}^{i\gamma_{\mathbf{k}}}\\ [-d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]\mathrm{e}^{-i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})\mathrm{e}^{-i\gamma_{\mathbf{k}}}&-\xi_{\mathbf{k}}+g|\mathbf{k}|&0\\ d_{3}(\mathbf{k})\mathrm{e}^{-i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]\mathrm{e}^{-i\gamma_{\mathbf{k}}}&0&-\xi_{\mathbf{k}}-g|\mathbf{k}|\end{pmatrix},$ (91)
@@
$=\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}&0&0\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&0&0\\ 0&0&\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ 0&0&\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\end{pmatrix}\cdot\frac{1}{2\sqrt{E_{\mathbf{k}\perp}}}$
@@
$\begin{pmatrix}\mathrm{e}^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\mathrm{e}^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-\mathrm{e}^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-\mathrm{e}^{i\gamma_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ \sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\perp}-g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\perp}+g|\mathbf{k}|)(E_{\mathbf{k}\perp}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\end{pmatrix},$
@@
$d_{12}(\mathbf{k})\mathrm{e}^{i\Omega_{\mathbf{k}}}$ $=d_{1}(\mathbf{k})+id_{2}(\mathbf{k}),$
@@
$u_{\mathbf{k}}$ $=\frac{\mathrm{e}^{i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}\cos\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}2+}b_{\mathbf{k}-}&\cos\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}+\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}1-}b_{\mathbf{k}+}\\ \mathrm{e}^{i\varphi_{\mathbf{k}}}(\sin\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}+\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}2+}b_{\mathbf{k}-})&\mathrm{e}^{i\varphi_{\mathbf{k}}}(\sin\frac{\theta_{\mathbf{k}}}{2}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos\frac{\theta_{\mathbf{k}}}{2}a_{\mathbf{k}1-}b_{\mathbf{k}+})\end{pmatrix},$ (98)
$v_{\mathbf{k}}$ $=\frac{1}{2}\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}+\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}&\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}\\ \mathrm{e}^{i\varphi_{\mathbf{k}}}(\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-})&\mathrm{e}^{i\varphi_{\mathbf{k}}}(\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}+\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+})\end{pmatrix},$